Problem 1. T he reaction of iron and water vapor results in an equilibrium 3 Fe(s) + 4H2O(g) Fe3O4(s) + 4H2(g) 근 and an equilibrium constant, K, of 4.0 at 850-c. what is the concentration of hydrogen present at equilibrium if the reaction is initiated with 31 g of H20 and excess Fe in a 13 liter vessel. Assume the volumes of the Fe and Fe304 are small compared to the size of the vessel.
Answer #1

concentration of water vapor = (31/18)*(1/13) = 0.132 M     


   3 Fe(s)   +   4H2O(g) <-----> Fe3O4(s) + 4H2(g)

initial   excess       0.132 M           0         0

change                  -4x                        4x

equilibrium             0.132-4x                   4x

        
             Kc = [H2]^4/[H2O]^4

              4 = (4X)^4/(0.132-4X)^4

             X = 0.0193

At equilibrium,

[H2] = 4X = 4*0.0193 = 0.0772 M

[H2O] = 0.132-4X = 0.132-4*0.0193 = 0.0548 M